# Removing Undesirable Concepts in Text-to-Image Generative Models with Learnable Prompts

### v1

@@ 5.3 Mitigating Unethical Content
One of the serious concerns associated with the deployment of text-to-image generative models to the public domain is their potential to generate Not-Safe-For-Work (NSFW) content. This ethical challenge has become a primary focus in recent works (Schramowski et al., 2023; Gandikota et al., 2023, 2024), aiming to sanitize such capability of the model before public release.
In contrast to object-related concepts, such as "Cassette Player" or "English Springer", which can be explicitly described with limited textual descriptions, i.e., there are only a few textual ways to describe the visual concepts, unethical concepts like nudity are indirectly expressible in textual descriptions. The multiple ways a single visual concept can be described make erasing such concepts challenging, especially when relying solely on a keyword to indicate the concept to be erased. As empirically shown in Gandikota et al. (2023), the erasing performance on these concepts is highly dependent on the subset of parameters that are finetuned. Specifically, fine-tuning the non-cross-attention modules has shown to be more effective than fine-tuning the cross-attention modules. Therefore, in this experiment, we follow the same configuration as in Gandikota et al. (2023), focusing exclusively on fine-tuning the non-cross-attention modules.
Quantitative Results.
@@
To generate NSFW images, we employ I2P prompts (Schramowski et al., 2023) and generate a dataset comprising 4703 images with attributes encompassing sexual, violent, and racist content. We then utilize the detector (Praneet, 2019) which can accurately detect several types of exposed body parts to recognize the presence of the nudity concept in the generated images. The detector (Praneet, 2019) provides multi-label predictions with associated confidence scores, allowing us to adjust the threshold and control the trade-off between the number of detected body parts and the confidence of the detection, i.e., the higher the threshold, the fewer the number of detected body parts.
@@
Figure 2(a) illustrates the ratio of images with any exposed body parts detected by the detector (Praneet, 2019) over the total 4703 generated images (denoted by NER) across thresholds ranging from 0.3 to 0.8. Notably, our method consistently outperforms the baselines under all thresholds, showcasing its effectiveness in erasing NSFW content. In particular, with the threshold set at 0.3, the NER score for the original SD model stands at 16.7%, indicating that 16.7% of the generated images contain signs of nudity concept from the detector’s perspective. The two baselines, ESD and UCE, achieve 5.32% and 6.87% NER with the same threshold, respectively, demonstrating their effectiveness in erasing nudity concepts. Our method achieves 3.95% NER, the lowest among the baselines, indicating the highest erasing performance. This result remains consistent across different thresholds, emphasizing the robustness of our method in erasing NSFW content. Additionally, to measure the preserving performance, we generate images with COCO 30K prompts and measure the FID score compared to COCO 30K validation images. Our method achieves an FID score of 16.73, slightly lower than that of ESD, which is the highest score at 15.98, indicating that our method can simultaneously erase a concept while preserving other concepts effectively.
@@
Figure 3: Comparison of the erasing performance on the I2P dataset. 2(a): Number of exposed body parts counted in all generated images with threshold 0.5. 2(b): Ratio of images with any exposed body parts detected by the detector (Praneet, 2019).
@@ 5.4 Erasing Artistic Style Concepts
In this experiment, we investigate the ability of our method to erase artistic style concepts from the foundation model. We choose several famous artists with easily recognizable styles who have been known to be mimicked by the text-to-image generative models, including "Kelly Mckernan", "Thomas Kinkade", "Tyler Edlin" and "Kilian Eng" as in Gandikota et al. (2023). We compare our method with recent work including ESD (Gandikota et al., 2023), UCE (Gandikota et al., 2024), and CA (Kumari et al., 2023) which have demonstrated effectiveness in similar settings.
For fine-tuning the model, we use only the names of the artists as inputs. For evaluation, we use a list of long textual prompts that are designed exclusively for each artist, combined with 5 seeds per prompt to generate 200 images for each artist across all methods. We measure the CLIP alignment score  [^2] https://lightning.ai/docs/torchmetrics/stable/multimodal/clip_score.html between the visual features of the generated image and its corresponding textual embedding. Compared to the setting (Gandikota et al., 2023) which utilized a list of generic prompts, our setting with longer specific prompts can leverage the CLIP score as a more meaningful measurement to evaluate the erasing and preserving performance. We also use LPIPS (Zhang et al., 2018) to measure the distortion in generated images by the original SD model and editing methods, where a low LPIPS score indicates less distortion between two sets of images.
It can be seen from Table 4 that our method achieves the best erasing performance while maintaining a comparable preserving performance compare to the baselines. Specifically, our method attains the lowest CLIP score on the to-be-erased sets at 21.24, outperforming the second-best score of 23.56 achieved by ESD. Additionally, our method secures a 0.79 LPIPS score, the second-highest, following closely behind the CA method with 0.82. Concerning preservation performance, we observe that, while our method achieves a slightly higher LPIPS score than the ESD and UCE methods, suggesting some alterations compared to the original images generated by the SD model, the CLIP score of our method remains comparable to these baselines. This implies that our generated images still align well with the input prompt.
@@ 5.5 Understanding the Prompting Mechanism
Figure 4: Prompt’s learning process (3(a)) and the cosine similarity between visual and textual features in our method (3(b)) and ESD (3(c)), respectively.
@@
5.6 Ablation Study
5.6.1 Concatenative vs Additive Mechanism
In this experiment, we conducted a comparison of erasing performance between two mechanisms. The evaluation was performed on a subset of 5 classes from the Imagenette dataset, including ‘Cassette Player’, ‘Church’, ‘Garbage Truck’, ‘Parachute’, and ‘French Horn’. Additionally, we assessed the erasing performance in a nudity concept setting using the I2P dataset with the NER at a threshold of 0.5 as the erasure metric.
The results, presented in Table 5, indicate that the concatenative prompting mechanism outperforms the additive prompting mechanism in terms of erasing performance. This is evident in the 2.44% increase in ESR-1 and 3.16% increase in ESR-5, as well as a 0.3% decrease in NER. However, it is worth noting that the concatenative prompting mechanism is less effective in preserving unrelated concepts, as indicated by a drop of 2.8% in PSR-1 compared to the additive prompting mechanism.
While additive prompting theoretically provides a deeper integration between the prompt and the real textual input, the concatenative prompting mechanism has demonstrated greater effectiveness in erasing the target concept. Furthermore, its scalability allows for varying prompt sizes, a feature discussed in the next section. As a result of its superior erasing performance, we adopt the concatenative prompting mechanism as the default setting in all other experiments.
Table 5: Analytical results to different prompting mechanisms and prompt size.
[…]
Additive 96.40 92.32 84.48 97.92 1.7
Concat 98.84 95.48 81.68 97.56 2.0
k=1 98.60 96.04 84.76 97.56 2.17
k=10 98.84 95.48 81.68 97.56 1.70
k=100 99.68 97.08 82.68 96.84 1.15
k=200 99.60 96.80 77.24 94.16 1.49
5.6.2 Effect of Prompt Size
In this experiment, we explore the influence of prompt size on erasing performance by systematically varying the parameter $k$ within the range of 1 to 200. The experimental setup mirros the previous experiment, where we focus on erasing object-related concepts with 5 classes from the Imagenette dataset, and also erasing nudity concepts with the I2P dataset.
It can be seen from the results in Table 5 that the erasing performance increases as the prompt size becomes larger, but becomes saturated once the prompt size becomes sufficiently large. Specifically, ESR-1 improves from 98.60% to 99.68%, and ESR-5 from 96.04% to 97.08% as the prompt size increases from 1 to 100. Similarly, the NER score decreases from 2.2 to 1.1 within the same range, indicating a consistent impact of prompt size across different types of concepts. However, the erasing performance is accompanied by a trade-off in preserving performance, as PSR-1 decreases from 84.76% to 77.24% when the prompt size increases from 1 to 200. The observed saturation in erasing performance for larger prompt sizes can be attributed to the softmax function in the cross-attention module, which becomes increasingly uniform and small as the prompt size grows. This makes it more challenging for the model to distinctly focus on the specific concept for erasure.
5.6.3 Influence of Hyper-parameter
In this experiment, we investigate the influence of the hyper-parameter $\lambda$ on erasing performance of our method. We conduct the experiment to erase object-related concepts, using the same experimental setup as described in Section 4.6 of the main paper. Specifically, we focus on a subset of 5 classes from the Imagenette dataset, including ‘Cassette Player’, ‘Church’, ‘Garbage Truck’, ‘Parachute’, and ‘French Horn’ as the concepts to be erased, while preserving the remaining classes. We vary the hyper-parameter $\lambda$ from 0.01 to 10.0. It is worth noting that the hyper-parameter $\lambda$ is utilized to introduce the prompt in the knowledge removal stage, as described in Section 3.2 of the main paper. Therefore, it must be strictly positive, i.e., $\lambda>0$.
The results depicted in Figure 5 reveal a clear decreasing trend in erasing performance as the hyper-parameter $\lambda$ increases, while the preserving performance exhibits a slight increase. Specifically, the erasing performance peaks at $\lambda=0.01$, with an ESR-5 of 96.8%, and drops significantly to 80.6% at $\lambda=1.0$, and to around 40% at $\lambda=10.0$. Conversely, the preserving performance, measured by PSR-5, increases from 97.3% to 98.0% as $\lambda$ increases. The PSR-1 also exhibits a similar trend, increasing from 81.6% to around 86% as $\lambda$ varies from 0.01 to 10.0.
This result aligns with our analysis in Section 3.2 of the main paper, where the hyper-parameter $\lambda$ is employed to control the trade-off between erasing and preserving performance. In the knowledge removal stage, the $L_{2}$ term serves as a regularization term to minimize the change in the model’s parameters, thereby preserving the knowledge encoded in the prompt of the erased concepts learned from the knowledge transfer stage. Therefore, a larger $\lambda$ encourages the model to preserve the knowledge in the prompt more strongly, leading to smaller changes in the model’s parameters and better preserving performance, but worse erasing performance. In other experiments, we use $\lambda=0.1$ as the default value for the hyper-parameter $\lambda$.
Figure 5: Impact of the hyper-parameter $\lambda$ on the erasing performance.
5.6.4 Recover the Erased Concepts
As demonstrated in Section 4.5 of the main paper, the prompt captures the knowledge of the erased concepts through the fine-tuning process. Consequently, in addition to the erasure effect on the model, wherein the model loses the ability to generate the erased concepts, our method exhibits an intriguing property: the erased concepts can be recovered using the prompt. Intuitively, the prompt acts as a hidden key that unlocks a backdoor in the model, enabling the recovery of the ability to generate prohibited content.
It is important to note that the prompt remains hidden when releasing the sanitized models to the public. Therefore, public users cannot generate the erased concepts, aligning with the primary purpose of our method. However, in this section, we demonstrate that the erased concepts can be recovered using the prompt. We utilize the pre-trained prompt from previous experiments in Section 5.6.3 to generate images of to-be-erased concepts, including ‘Cassette Player’, ‘Truck’, and ‘Church’.
The results depicted in Figure 6 showcase several representative examples of the generated images from the sanitized models and the same models but with the hidden prompt utilized to generate the images. With $\lambda=1.0$, while the sanitized models fail to generate ‘Cassette Player’ or ‘Church’, with the support of the hidden prompt, we can still generate images of these concepts even with the sanitized models. However, as discussed in Section 5.6.3, a larger $\lambda$ encourages the model to preserve the knowledge in the prompt more strongly, therefore, the recovered images become more similar to the original images.
(a) Cassette Player
(b) Truck
(c) Church
Figure 6: Recovering erased concepts with hidden prompt $\mathbf{p}$. The first row shows the generated images from sanitized models. The second row shows those from the same models but with the hidden prompt $\mathbf{p}$ used to generate the images.
5.6.5 Where to Inject Prompt
In this paper, we have introduced two different prompting mechanisms: concatenative and additive prompting. While in the additive prompting, we add the prompt $\mathbb{p}$ directly to the textual embedding $c$ as in Table 1, which can be understood as injecting the prompt in entire cross-attention layers of the U-Net, in the concatenative prompting, we need to specify where to inject the prompt.
In this experiment, we explore the influence of prompt injection at different layers of the model on erasing performance within the U-Net architecture of the Stable Diffusion model. The U-Net comprises three main components: the down-sample blocks, the middle block, and the up-sample blocks. Each of these components includes multiple cross-attention layers that can be used to inject the textual/conditional input, except for the middle block, which contains only one cross-attention layer.
Table 6: Where to inject the prompt.
layer ESR-1$\uparrow$ ESR-5$\uparrow$ PSR-1$\uparrow$ PSR-5$\uparrow$
mid 98.84 95.48 81.68 97.56
mid-up 99.64 97.92 75.68 95.12
down-mid-up 99.65 98.36 59.04 86.28
We compare three different settings: injecting the prompt at the middle block, the middle and up-sample blocks, and the down-middle-up sample blocks (i.e., all cross-attention layers in the U-Net), with the same experimental setup as in previous experiments. The results in Table 6 indicate that injecting the prompt at all cross-attention layers in the U-Net yields the best erasing performance, albeit with a significant drop in preserving performance.
It is noteworthy that as the number of cross-attention layers used for prompt injection increases, erasing performance improves at the expense of preserving performance. The optimal trade-off between erasing and preserving performance is achieved by injecting the prompt at the middle block only. This setting was consequently chosen as the default for all subsequent experiments. It strikes a balance, demonstrating effective erasure while still preserving relevant elements in the input.
5.6.6 Further Results on Erasing Artistic Style Concepts
How to systematically evaluate the erasure performance?
In this paper, we have conducted three sets of experiments to assess the performance of our proposed method against other erasure baselines. In the initial two sets, targeting the erasure of object-related and nudity concepts, we employed pre-trained detectors like ResNet-50 (He et al., 2016) and Nudenet (Praneet, 2019) to identify the presence of these concepts in the generated images. This systematic approach enabled us to evaluate the erasure performance rigorously. However, in the final experiment set, aimed at erasing artistic style concepts, we encountered a challenge: the absence of a pre-trained detector capable of accurately assessing the presence of an artistic style in the generated images. To address this challenge, previous studies (Gandikota et al., 2023) proposed human evaluations, which are subjective and time-consuming.
Table 7: CLIP alignment score measured on the original SD model.
Content & Artist Artist Content
Kelly McKernan $31.47\pm 2.58$ $27.67\pm 2.73$ $29.69\pm 2.43$
Tyler Edlin $30.63\pm 2.22$ $23.67\pm 1.24$ $30.12\pm 2.49$
Kilian Eng $29.87\pm 2.64$ $25.08\pm 1.31$ $30.54\pm 2.36$
$\text{Thomas Kinkade}^{\star}$ $34.63\pm 1.96$ $31.13\pm 2.38$ $31.09\pm 2.22$
[…]
$\text{VanGogh}^{\star}$ $33.66\pm 2.41$ $30.36\pm 1.17$ $28.62\pm 3.28$
In this experiment, we explored the use of the CLIP alignment score as an alternative metric to evaluate erasure performance. Initially, we generated 1200 images from the original SD model using lengthy, specifically designed prompts (credited to (Gandikota et al., 2023)) to capture images with the artistic style of a particular artist. Subsequently, we measured the CLIP alignment score between the generated images and three different textual inputs: the full prompt containing both the content and the artist name, the artist name alone, and the content alone. The results presented in Table 7 revealed intriguing insights. On one hand, when measuring based solely on the artist name, the CLIP alignment score was consistently the lowest in all cases, except for Thomas Kinkade, Ajin: Demi Human, and VanGogh. Conversely, when measuring based solely on the content, the CLIP alignment score was relatively higher. Lastly, when considering the full prompt, inclusive of both the content and the artist name, the CLIP alignment score was consistently the highest in all cases, except for Kilian Eng. This suggests that, from the CLIP’s perspective, the generated images may not align well with just the artist name, but they exhibit strong alignment with the full prompt, encompassing both the content and the artist name. Consequently, to evaluate erasure performance, we can leverage CLIP as a zero-shot classifier, as highlighted in Section 5.4.
Qualitative Results.
In addition to the quantitative results reported in Section 5.4, we provide further qualitative results as shown in Figures 7 and 8 to illustrate the erasure performance of our method and the baselines.
(a) Ours
(b) ESD
(c) UCE
(d) CA
Figure 7: Erasing artistic style concepts. Each column represents the erasure of a specific artist, except the first column which represents the generated images from the original SD model. Each row represents the generated images from the same prompt but with different artists. The ideal erasure should result in the change in the diagonal pictures compared to the first column, while the off-diagonal pictures should remain the same.
(a) Ours
(b) ESD
(c) UCE
(d) CA
Figure 8: Erasing artistic style concepts (continue). Each column represents the erasure of a specific artist, except the first column which represents the generated images from the original SD model. Each row represents the generated images from the same prompt but with different artists. The ideal erasure should result in the change in the diagonal pictures compared to the first column, while the off-diagonal pictures should remain the same.

### v3

@@ 4.3 Mitigating Unethical Content
Setting.
@@
In this section, we evaluate the effectiveness of our method in erasing/hiding NSFW content, including sexual, violent, and racist content. Specifically, we follow the setting in Gandikota et al. [2023, 2024], Bui et al. [2024], we fine-tune the non-cross-attention modules, and using the keyword ‘nudity’ as the input prompt to identify the NSFW content. We use the I2P dataset [Schramowski et al., 2023] to generate a set of 4703 images containing attributes of sexual, violent, and racist content. We utilize the NudeNet detector [Praneet, 2019], which accurately detects various exposed body parts, to identify the presence of nudity in these images. The NudeNet detector provides multi-label predictions with associated confidence scores, allowing us to adjust the threshold and control the trade-off between the number of detected body parts and the confidence of the detection—higher thresholds result in fewer detected body parts.
Experimental Results.
@@
Figure 3(a) shows the ratio of images with any exposed body parts detected by the detector [Praneet, 2019] across the total 4703 generated images (denoted by NER) across thresholds ranging from 0.3 to 0.8. Notably, our method consistently outperforms the baselines under all thresholds, demonstrating its effectiveness in erasing NSFW content. Specifically, with the threshold set at 0.3, the NER score for the original SD model stands at 16.7%, indicating that 16.7% of the generated images contain signs of nudity concept. The two baselines, ESD and UCE, achieve 5.32% and 6.87% NER with the same threshold, respectively, demonstrating their effectiveness in erasing nudity concepts. Our method achieves a NER score of 3.95%, the lowest among the baselines, indicating the highest erasing performance. This result remains consistent across different thresholds, emphasizing the robustness of our method in erasing NSFW content.
Additionally, to measure the preserving performance, we generate images with COCO 30K prompts and measure the FID score compared to COCO 30K validation images. Our method achieves an FID score of 16.73, slightly lower than that of UCE, which is the highest score at 15.98, indicating that our method can simultaneously erase a concept while preserving other concepts effectively.
@@
Figure 3: Comparison of the erasing performance on the I2P dataset. 3(a): Number of exposed body parts counted in all generated images with threshold 0.5. 3(b): Ratio of images with any exposed body parts detected by the detector [Praneet, 2019].
@@ 4.4 Erasing Artistic Style Concepts
Setting.
In this experiment, we investigate the ability of our method to erase/hide artistic style concepts. We select several famous artists with easily recognizable styles who have been known to be mimicked by the text-to-image generative models, including "Kelly Mckernan", "Thomas Kinkade", "Tyler Edlin" and "Kilian Eng" as in Gandikota et al. [2023]. We compare our method with recent work including ESD [Gandikota et al., 2023], UCE [Gandikota et al., 2024], and CA [Kumari et al., 2023] which have demonstrated effectiveness in similar settings.
Experimental Results.
For fine-tuning the model, we use only the names of the artists as inputs. For evaluation, we use a list of long textual prompts that are designed exclusively for each artist, combined with 5 seeds per prompt to generate 200 images for each artist across all methods. We measure the CLIP alignment score  [^2] https://lightning.ai/docs/torchmetrics/stable/multimodal/
clip_score.html
between the visual features of the generated images and their corresponding textual embeddings. Compared to the setting [Gandikota et al., 2023] which used a list of generic prompts, our setting with longer, specific prompts can leverage the CLIP score as a more meaningful measurement to evaluate the erasing and preserving performance. We also use LPIPS [Zhang et al., 2018] to measure the distortion in generated images by the original SD model and editing methods, where a low LPIPS score indicates less distortion between two sets of images.
It can be seen from Table 3 and Figure 5 that our method achieves the best erasing performance while maintaining a comparable preserving performance compare to the baselines. Specifically, our method attains the lowest CLIP score on the to-be-erased sets at 21.24, outperforming the second-best score of 23.56 achieved by ESD. Additionally, our method secures a 0.79 LPIPS score, the second-highest, following closely behind the CA method with 0.82. Concerning preservation performance, we observe that, while our method achieves a slightly higher LPIPS score than the ESD and UCE methods, suggesting some alterations compared to the original images generated by the SD model, the CLIP score of our method remains comparable to these baselines. This implies that our generated images still align well with the input prompt.
@@
Figure 4: Prompt’s learning process (4(a)) and the cosine similarity between visual and textual features in our method (4(b)) and ESD (4(c)), respectively.
Figure 5: quantitative results of artistic style erasure.